# Diffusion Posterior Sampling is Computationally Intractable

### v1

@@ Appendix F Well-Modeled Distributions Have Accurate Unconditional Samplers
Theorem F.2 (Unconditional Sampling Theorem, Implied by [BBDD24], adapted from [GPPX23]).
@@
$\operatorname{\mathbb{P}}\left[\|\widehat{x}-x\|>\frac{1}{d^{C/6}}\right]<\frac{1}{d^{C/6}}$
@@ Appendix G Cryptographic Hardness
A one-way function $f$ is a function such that every polynomial-time algorithm fails to find a pre-image of a random output of $f$ with high probability.
Definition G.1.
A function $f:\{\pm 1\}^{n}\to\{\pm 1\}^{m(n)}$ is one-way if, for any polynomial-time algorithm $\mathcal{A}$, any constant $c>0$, and all sufficiently large $n$,
$\operatorname{\mathbb{P}}_{x\sim\mathcal{U}_{n}}\left[f(\mathcal{A}(f(x)))=f(x)\right]<n^{-c}.$
@@
For $l(n)>m(n)$, we just need to pad $l(n)-m(n)$ 1’s at the end of the output, i.e.,
@@
In the realm of $\{+1,-1\}$, $-1$ corresponds to True and $+1$ corresponds to False. We can use a layer of neurons to translate it to $\{0,1\}$ first, where $1$ corresponds to True and $-1$ corresponds to False. We will translate $\{0,1\}$ back to $\{+1,-1\}$ when output.
@@ Appendix H Utility Results
$\operatorname{\mathbb{P}}\left[\sup_{y\in[x-\gamma,x+\gamma]}s(y)\geq t\right]\leq e^{-\sigma t}$
@@
$\operatorname{\mathbb{P}}\left[\sup_{y\in[x-\gamma,x+\gamma]}s(x)^{k}\geq t^{k}\right]\leq\frac{\operatorname*{\mathbb{E}}\left[\sup_{y\in[x-\gamma,x+\gamma]}s(x)^{k}\right]}{t^{k}}\leq\left(\frac{15k}{t\sigma}\right)^{k}$
@@
$\operatorname{\mathbb{P}}\left[\sup_{y\in[x-\gamma,x+\gamma]}|s(x)|\geq\frac{15}{e\sigma}\log\frac{1}{\delta}\right]\leq\delta$
@@
Lemma H.3 (Lemma A.1 from [GLPV22]).
@@
This proof is given in Lemma A.2 of [GLPV22], and is reproduced here for convenience and completeness, since a statement in the middle of their proof is what we use.
@@
Lemma H.5 (Lemma 3.1 from [GLPV22]).
@@
Lemma H.6 (Lemma B.3 from [GLPV22]).
@@
Lemma H.7 (Lemma A.6 from [GLPV22]).
@@
Lemma H.9 (Laurent-Massart Bounds[LM00]).
@@
$\operatorname{\mathbb{P}}[\left\|v\right\|^{2}-n\geq 2\sqrt{nt}+2t]\leq e^{-t}.$
Lemma H.10 (See [MRT18, Lemma 6.27]).

### v2

@@ Appendix F Well-Modeled Distributions Have Accurate Unconditional Samplers
Theorem F.2 (Unconditional Sampling Theorem, Implied by [benton2024nearly], adapted from [gupta2023sampleefficient]).
@@
$\operatorname*{\mathbf{Pr}}\left[\|\widehat{x}-x\|>\frac{1}{d^{C/6}}\right]<\frac{1}{d^{C/6}}$
@@ Appendix G Cryptographic Hardness
Recall that a one-way function $f$ is a function such that every polynomial-time algorithm fails to find a pre-image of a random output of $f$ with high probability.
@@
For $l(n)>m(n)$, we just need to pad l(n)m(n)l(n)-m(n) 1’s at the end of the output, i.e.,
@@
In the realm of $\{+1,-1\}$, 1-1 corresponds to True and +1+1 corresponds to False. We can use a layer of neurons to translate it to $\{0,1\}$ first, where $1$ corresponds to True and 1-1 corresponds to False. We will translate $\{0,1\}$ back to $\{+1,-1\}$ when output.
@@ Appendix H Utility Results
$\operatorname*{\mathbf{Pr}}\left[\sup_{y\in[x-\gamma,x+\gamma]}s(y)\geq t\right]\leq e^{-\sigma t}$
@@
$\operatorname*{\mathbf{Pr}}\left[\sup_{y\in[x-\gamma,x+\gamma]}s(x)^{k}\geq t^{k}\right]\leq\frac{\operatorname*{\mathbb{E}}\left[\sup_{y\in[x-\gamma,x+\gamma]}s(x)^{k}\right]}{t^{k}}\leq\left(\frac{15k}{t\sigma}\right)^{k}$
@@
$\operatorname*{\mathbf{Pr}}\left[\sup_{y\in[x-\gamma,x+\gamma]}|s(x)|\geq\frac{15}{e\sigma}\log\frac{1}{\delta}\right]\leq\delta$
@@
Lemma H.3 (Lemma A.1 from [gupta2022finitesample]).
@@
This proof is given in Lemma A.2 of [gupta2022finitesample], and is reproduced here for convenience and completeness, since a statement in the middle of their proof is what we use.
@@
Lemma H.5 (Lemma 3.1 from [gupta2022finitesample]).
@@
Lemma H.6 (Lemma B.3 from [gupta2022finitesample]).
@@
Lemma H.7 (Lemma A.6 from [gupta2022finitesample]).
@@
Lemma H.9 (Laurent-Massart Bounds[LaurentMassart]).
@@
$\operatorname*{\mathbf{Pr}}[\left\|v\right\|^{2}-n\geq 2\sqrt{nt}+2t]\leq e^{-t}.$
Lemma H.10 (See mohri2018foundations, Lemma 6.27).